OpenStudy (anonymous):

change point into rectangular for (4,4pi/3) need help don't understand how to do this type of questions. A.(-2,sqrt3) B. (-2,2sqrt3) C.(2, -2sqrt3) D. none

OpenStudy (ddcamp):

The point your given is in polar form. The first number is the radius (r), the second is the angle from the positive x-axis (theta).: 

OpenStudy (ddcamp):

using trig, we can find the formulas: x= r*cos(theta) y = r*sin(theta) (x,y) is the point in rectangular coordinates

OpenStudy (jdoe0001):

\(\large \left(\begin{array}{cccllll} &4,&\cfrac{4\pi}{3}\\ \hline\\ &\uparrow &\uparrow \\ &r&\theta \end{array}\right)\qquad \begin{array}{llll} x=rcos(\theta)\\ y=rsin(\theta) \end{array}\)

OpenStudy (jdoe0001):

so just find "x" and "y" and your rectangular coordinates will be (x,y) :) as DDCamp said